eric is ten years younger than leslie. in ten years, he will be ten years more than one-half as old as leslie. find their present age.
Answer by htmentor(1343)

eric is ten years younger than leslie. in ten years, he will be ten years more than one-half as old as leslie. find their present age.
==============================================================
Let L = Leslie's present age
Then Eric's age = L - 10
In 10 years Leslie's age will be L + 10 and Eric's age will be (L-10) + 10 = L
Since in 10 years Eric will be 10 more then 1/2 of Leslie, we can write
L = 1/2(L+10) + 10
Solve for L:
L = L/2 + 15
L = 30
Leslie's age = 30
Eric's age = 20